If A and B are independent events, P(A) = 0.15, and P(B) = 0.75, find the probabilities below. (Enter your answers to four decimal places.)
(a) P(A ∩ B)

(b) P(A ∪ B)

(c) P(A | B)

(d) P(Ac ∪ Bc)


Answer by greenestamps(13195)



(a) (.15)*(.75) = .1125

(b) P(A) + P(B) - P(A and B) = .15 + .75 - .1125 = .7875

(c) P(A and B) / P(B) = .1125/.75 = .15

(d) 1 - P(A and B) = 1 - .1125 = .8875